Solve (matrix)
[-6, 3t+9]
[-1, 4]
=
[-6, 6t-4]
[-1, 3y+2]
for each variable.
1
Expert's answer
2012-02-09T08:24:45-0500
Two matrices are equal if and only if their elements are equal. So, we get the
system of 2 equations:
3t+9 = 6t-4
4 = 3y+2.
From the first equation we
have 3t = 13, t = 13/3; from the second 3y = 2, y = 2/3.
The solution: t =
13/3, y = 2/3.